Recurring PLL within a time

Totally, like, sometimes I get 4-5 PLL skips in a row.
Other times I get like, 20 PLL skips in a row.

yeah, the probability is just 1/14016833953562607293918185758734155776 .

You should play to some lotery.

(ZB method)

Yes, but we are using terminology incorrectly here ;)

PLL skip implies that you "skipped" the step as per chance. As in you had some chance to skip this step, and you got lucky and this occurred.

The chance of a "PLL" skip for ZBLL is actually the chance of a full LL skip. The reason is that PLL is not a step in the ZB method but a special case of the ZB method.

For example, Chris has a 1/27 chance to have a fully oriented last layer after solving ZBF2L. This means that Chris will do a standard PLL alg when this happens, most of the time. Now, given that Chris has ended up with an oriented last layer after ZBF2L (1/27 chance to do so), he then has a 1/72 chance to skip the resulting PLL case.

So Chris, the chance that you have a PLL skip is 1/27 * 1/72 = 1/1944 This of course assumes you don't do anything extra crazy to partially influence the orientation of corners during ZBF2L.

Remember that PLL is not a step in your method, it is a special sub-case of your method, the one that occurs 1/27 of the time when the last layer is oriented after ZBF2L. Using terminology in this way can lead to confusing claims of "I skipped PLL 20 times in a row" and counter-claims of "This is so completely improbable as to be effectively impossible".

;)

Chris
 
This only happens to me when I do solves with random scrambles.

random scrambles?
I mean non computer generated scrambles.


I don't understand how you could get one PLL more than others simply by not using a computer generated scramble. Unless you were doing the same scramble over and over again, I am certain that doing the cross, inserting 4 F2L pairs and doing your OLL will well and truly scramble the last layer.

Think of the "butterfly effect". If you do two scrambles by hand which are similar except for a small difference, after all the steps, that small difference will be amplified.

I also think that I get the same PLL many times in a row but I think that is just psychology. Many combinations of PLLs will occur but we just notice them more if there are chains of consecutive ones.
 
Yes, but we are using terminology incorrectly here
No, we're not.
By the way, I've gotten 100 consecutive solves with 1LLL. ;)

Whoa, ok now I am totally confused. :confused:

How is using a ZBLL alg considered at the same time a PLL skip? I would think it better to say that a property of ZBLL is that each alg permutes all pieces to their correct location. To me, "skip" implies there is some element of chance involved with the step. Sure you can improve your chances of a skip, but there is still some "luck" element as to whether or not the skip happens.

Chris would always permute all pieces using a ZBLL alg, so there is no luck element here. He has a 100% chance to not only orient the corners, but permute all pieces to their correct location. I don't see how this counts as a PLL "skip". I would interpret this by saying that the permutation of all pieces is handled during the same step as the orientation of corners. Thus PLL (implying all pieces are oriented correctly, but need to be permuted) only happens in the same case for Chris that he has a 1/27 chance of all corners being oriented after ZBF2L. Given that he has all oriented corners, he has the same chance, 1/72 to skip PLL. This results in a 1/1944 chance of a "PLL skip" and here I mean "skip" in the same sense of implying that luck was involved in the skipping of the step.

Chris
 
Whoa, ok now I am totally confused. :confused:

Chris

Lucas' LL consists of 'PLL', his f2l consists of 'Cross -> 3 pairs -> EO & Last f2l Edge -> CO & Last f2l Corner' :P

I know how MGLS works, and that Lucas uses it, so I can see how you can get huge numbers of solves with a 1LLL. Every solve with MGLS is a 1LLL thinking of it this way. That makes sense to me.

What confuses me is that Lucas says that we are not using terminology incorrectly by referring to a ZBLL as a PLL skip. That's what confuses me. Perhaps we have different defintions of what PLL is?

Chris
 
This only happens to me when I do solves with random scrambles.

random scrambles?
I mean non computer generated scrambles.


I don't understand how you could get one PLL more than others simply by not using a computer generated scramble. Unless you were doing the same scramble over and over again, I am certain that doing the cross, inserting 4 F2L pairs and doing your OLL will well and truly scramble the last layer.

Think of the "butterfly effect". If you do two scrambles by hand which are similar except for a small difference, after all the steps, that small difference will be amplified.

I also think that I get the same PLL many times in a row but I think that is just psychology. Many combinations of PLLs will occur but we just notice them more if there are chains of consecutive ones.

I always get the G perms I don't know, J's, and alot others. I rarely can dive right into PLL because the first case I get is one I don't know, which is why I notice how many times it comes up.
 
If PLL isn't a step in the method, it's jsut stupid to pretend to have a PLL skip.

The only reason I can see is to make the attention whore by saying « hey guys, sometinme I got 20 PLL skip in a row » and get amazed answers about you, without any interest.

Making people speak of you, that's the only point of this knid of affirmation.
 
...
I always get the G perms I don't know, J's, and alot others. I rarely can dive right into PLL because the first case I get is one I don't know, which is why I notice how many times it comes up.
(match the formatting)
And that is why I ask people for basic statistic knowledge.
So basically you get a lot of different PLL's but you think you always get the same one?
 
Back
Top